

********************************************************************
UARS WEEKLY STATUS REPORT - 15 OCT 2004
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully gathered daily mid (1.1 nm) resolution, MgII,
occultation, solar pointing offset, and continuous monitoring
scans for periods of UARS telemetry and tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004159 has been reprocessed and archived at the GSFC DAAC. HEPSA Level 2
data files from 2004160 to 2004179 are processed and waiting for archiving.
MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 2004139.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing recieved.
***********
FOT
***********
REPORTING PERIOD
This report covers 02 October 2004 (Orbit 71466, GMT Day 276) through
08 October 2004 (Orbit 71570, GMT Day 282).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 281/1730 71552 POWER OFF
The beta angle ranged between the angles of 30.0 degrees to 55.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 65 degrees on 13 October 2004 (DOY 287).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.58 to 1.90 degrees C. Additional
battery performance data for each battery on 01 October and 08 October
2004 are provided in Appendix D.
The clock error ranged from 11.0 to -10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
A problem was found last week in the new release of the CMS code. The
issue is believed to be with the namelist files, but software
development is looking into the exact cause. FDF has also been asked
to provide a User Antenna Views file, which contains SMA codes. This
is needed to complete the CMS processing of SMA supports. A meeting
will be held with FDF to discuss this issue. These two problems have
delayed the release of the new CMS code to operations. No new date
has been set for completion.
MEETINGS
The bi-weekly meeting with the HRDI PI team was held on 06 October
2004. At this meeting the FOT was requested to gather sun data prior
to the FHST-2 failure.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 November 2004 (DOY 307).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 October 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
71477 276/1814 04-284 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71494 277/2137 04-285 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71552 281/1730 04-286 HALOE POWER OFF
71556 282/0012 04-288 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71556 282/0012 04-287 MPS VT TO LEVEL-5
71558 282/0250 04-287 MPS BATSOC
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 October 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 October 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
71477 276/1814 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
71494 277/2131 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
71556 282/0039 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 October 2004
01 October 2004 - GMT Day 275 Beta = 28.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-3.68 +00.0/+00.0 07.0 0.00 0.0
2 00.0/-44.8 +5.79/+5.15 +27.2/-13.6 26.9 1.02 19.7
3 -16.8/-78.4 +4.51/+3.87 +26.0/-13.2 26.7 1.02 19.0
08 October 2004 - GMT Day 282 Beta = 55.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 -22.4/- 89.6 +4.50/+2.60 +23.2/-15.2 26.4 1.04 16.5
3 -22.4/-168.0 +3.23/+1.65 +22.0/-14.4 26.2 1.04 16.0
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
71556 282/0012 LV6 -> LV5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 October 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT 71501 278/0826 30832 Generic Late Acquisition
04-067 FOT 71552 281/1730 37177 Data Drop Out
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-063 FOT See List Sept 04 30832 Generic Late Acquisition
04-064 FOT 71435 273/2302 37111 Data Dropout / bad weather
04-065 FOT 71452 275/0256 37125 Bad Weather at GUAM
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-066 FOT See List Oct 04 30832 Generic Late Acquisition
04-067 FOT 71552 281/1730 37177 Data Drop Out